yeetantor:

A printer printed 68 letters in 4 seconds. If it prints letters at a constant speed, it can print ______ letters in 30 seconds.

jimthompson5910:

A printer printed 68 letters in 4 seconds How many letters can it print in 1 second?

yeetantor:

well it can print in 1 second that would be im not sure actually

jimthompson5910:

Think of it like an equation 68 letters = 4 seconds We can divide both sides by 4 to get 68/4 letters = 4/4 seconds 17 letters = 1 second This tells us the speed of the printer is 17 letters per second. Agreed?

yeetantor:

yea it seems like a answer to role with agreed

jimthompson5910:

So from here we multiply both sides by 30 to change the "1 second" into "30 seconds" 17 letters = 1 second 30*17 letters = 30*1 second 510 letters = 30 seconds The printer can print 510 letters in 30 seconds

yeetantor:

oh ok thanks dude big help i've been stuck on this question for about two hours

jimthompson5910:

If you're curious, here's another way to do the problem (number of letters)/(number of seconds) = (number of letters)/(number of seconds) (68 letters)/(4 seconds) = (x letters)/(30 seconds) 68/4 = x/30 68*30 = 4x ... cross multiply 2040 = 4x 4x = 2040 x = 2040/4 .... divide both sides by 4 x = 510 We get the same answer as before